OpenStudy (anonymous):

Help with.... The distance, d, an object travels is represented by the formula d = rt, where r is the rate at which the object is traveling and t is the duration of time over which the object travels. If d has units of meters and r has units of meters per minute, what must be the units of t? A. seconds B. meters C. meters per second D. minutes

OpenStudy (anonymous):

Plug in what you know for your values. It is given that d is in meters and that r is in meters/minute and that it is multiplied by "t" so what must t be?

OpenStudy (anonymous):

C.meters per second

OpenStudy (anonymous):

or A.seconds

OpenStudy (anonymous):

Why would you say that?

OpenStudy (anonymous):

because its meters then minutes so wouldnt it be seconds after

OpenStudy (anonymous):

It is not asking what would be after? It is telling you that after is meters and that before you have meters/minute multiplied by something to get meters only

OpenStudy (anonymous):

\[meters = \frac{ meters }{ minute } \times t\] What must t be so that you only have meters in your answer?

OpenStudy (anonymous):

meters

OpenStudy (anonymous):

Lets try it out...\[d = \frac{ meters }{ minute } \times t\] if we say t is equal to meters then we have \[d = \frac{ meters }{ minute } \times meters\] and then d would be\[d = \frac{ meters^2 }{ minute }\]

OpenStudy (anonymous):

We can obviously see that is not what they said d would be

OpenStudy (anonymous):

So which should be the answer so that you are only left with meters as your result?